A person of mass 70 kg stands at the center of a rotating merry-go-round platform of radius 3.4 m and moment of inertia 940 kg⋅m2 . The platform rotates without friction with angular velocity 1.6 rad/s . The person walks radially to the edge of the platform.

Calculate the angular velocity when the person reaches the edge. In rad/sec

Calculate the rotational kinetic energy of the system of platform plus person before and after the person's walk. In J.
Answer #1

a) The moment of inertia of the person is, person = mR2 Here m is the mass of the person, and R is the radius of the platfom.

b) The rotational kinetic energy of the system of platform plus person before persons walk ー(940kg·㎡)(1.60 rad/s)2 -1203.2J
Answer #2

SOLUTION  :


a.


M. I. of  platform = 940 kg-m^2


M. I. of the person when he stands at the centre initially = m r^2 = 70 * 0^2 = 0  kg0-m^2


M. I. of the person when he stands at the edge finally = m r^2 = 70 * 3.4^2 = 809.20 kg-m^2 


Angular velocity of platform initially = 1.6 rad/s 


Angular momentum remains constant whether the person is at centre or at the edge.


So,


Angular momentum initially = Angular momentum finally

M. I. Initially * angular velocity initially = M. I. finally * angular velocity finally

=> (940 + 0) * 1.6 = (940 + 809.20) * angular velocity finally

=> angular velocity finally = (940 * 1.6) / (1749.20) 

=> angular velocity finally = 0.8598 = 0.86  rad/s (ANSWER) 


b.


Rotational K.E. before person walks to the edge

= 1/2 M.I. initially * (angular velocity initially)^2

= 1/2 * 940 * 1.6^2 

= 1203.2 J (ANSWER).


Rotational K.E. after person walks and reaches the edge

= 1/2 M.I. finally * (angular velocity finally)^2

= 1/2 * 1749.20 * 0.8598 

= 751.98 J  (ANSWER).

answered by: Tulsiram Garg